OpenStudy (anonymous):

When finding the sum of the arithmetic series 4 + 1 + -2 + -5 + -8 + -11, you will have a step of work that contains the number -3. Use complete sentences to describe how the -3 was calculated and what it means.

OpenStudy (texaschic101):

-3 is the common difference in this series. To find the common difference, take the second term (1) and subtract the first term (4). 1 - 4 = -3 ....you will find that if you add -3 to each term, you will get the next term. 4 + (-3) = 4 - 3 = 1 1 + (-3) = 1 - 3 = -2 -2 + (-3) = -2 - 3 = -5 etc...